OpenStudy (anonymous):

-2y - x = -2 -x - 4y = -6 elimination method please help no clue still

OpenStudy (anonymous):

In this case we multiply the 2nd by -1 and add it to the first one. So we get this: 4y-2y=6-2 2y=4, y=2 Now we go back to the first equation and we get this: -2*2-x=-2 x=2-4=-2 So the answer is: x=-2, y=2

OpenStudy (anonymous):

ur a life saver:)

OpenStudy (anonymous):

Thanks you! Does this makes sense? Do you think that you can do a similar one by your own?

OpenStudy (anonymous):

the ones i am given for my work arent simple :( see 2x + 10y = -42 -3x + 9y = -81